Find the length of the arc of the parabola y2 = x from (0, 0) to (1, 1)
1
Expert's answer
2020-06-30T17:25:20-0400

We have an explicitly defined function "f(y) = y^2".

So the length of this curve on the segment "y \in [0, 1]" will be calculated using the formula:


"L = \intop_0^1\sqrt{1 + {f'(y)}^2}dy" , then "f'(y) = 2y" "\implies L = \intop_0^1\sqrt{1 + 4y^2}dy".



"\intop_0^1\sqrt{1 + 4y^2}dy ="


"= \begin{bmatrix}
 u=\sqrt{1 + 4y^2} & du=\frac{4y}{\sqrt{1 + 4y^2} }\\
 dv = dy & d = y
\end{bmatrix}
= y\sqrt{1+4y^2}\mid_0^1-"


"- \int_0^1\frac{4y^2}{\sqrt{1+4y^2}}dy" .


"\int_0^1\frac{4y^2}{\sqrt{1+4y^2}}dy = \int_0^1\frac{1+4y^2-1}{\sqrt{1+4y^2}}dy =\int_0^1\sqrt{1+4y^2}dy -"


"-\int_0^1\frac{dy}{\sqrt{1+4y^2}}" .


So, "L= y\sqrt{1+4y^2}\mid_0^1-L +\int_0^1\frac{dy}{\sqrt{1+4y^2}}".


"2L = y\sqrt{1+4y^2}\mid_0^1+\int_0^1\frac{dy}{\sqrt{1+4y^2}}".


"L = \frac{\sqrt{5}}{2} + \frac{1}{2}\times\frac{1}{2}\int_0^1\frac{d(2y)}{\sqrt{1+(2y)^2}}".


"L = \frac{\sqrt{5}}{2} + \frac{1}{4}ln\mid2y + \sqrt{1+4y^2}\mid\vert_0^1" .


"L = \frac{\sqrt{5}}{2} + \frac{1}{4}ln\mid2+\sqrt{5}\mid - \frac{1}{4}ln\mid1\mid".


But "ln\mid1\mid = 0", then


"L = \frac{\sqrt{5}}{2} + \frac{1}{4}ln\mid2+\sqrt{5}\mid \approx1.4789".


Answer: 1.4789